1/1

Михаил Юрьевич Лермонтов, Михаил Юрьевич Лермонтов,

Михаил Юрьевич Лермонтов,

Парус  Das Segel  Sail
tradukita de Heinrich Greif  tradukita de A. Z. Foreman

Белеет парус одинокий В тумане моря голубом!.. Что ищет он в стране

далекой? Что кинул он в краю

родном?..

Wo Meer und Himmel sich ver-

einen, Erglänzt ein Segel, weiß und

weit - Was trieb es aus dem Land der

Seinen? Was sucht es in der Einsamkeit?

A sail drifts white and on its own Amid the light blue ocean haze. What does it seek in distant coun-

try? What made it leave its native

bays?

Играют волны, ветер

свищет, И мачта гнётся и

скрипит, Увы! – он счастия не

ищет, И не от счастия бежит! – Es pfeift der Wind. Die Wellen

drohen. Es knarrt der Mast. Das Segel

schwebt Nicht vor dem Glück ist es ge-

flohen. Es ist nicht Glück, wonach es

strebt.

Great billows play. High winds are

whistling Down at the bending, creaking

mast Oh! This one seeks no happy en-

ding And does not flee a happy past.

Под ним струя светлей

лазуры, Над ним луч солнца

золотой: – А он мятежный, просит

бури, Как будто в бурях есть

покой.

Strahlt auch in Gold der Him-

melsbogen, Und glänzt auch noch so blau

das Meer - Das Segel lechzt nach Sturm

und Wogen, Als ob in Stürmen Ruhe wär.

Beneath, a brighter stream than

azure. Above, the golden sunray flows Yet this one, restive...quests for

tempests As if in tempests were repose.

Verkinto de tiu ĉi Rusa poemo estas Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27).

Arg-545-1093 (2010-06-20 16:02:28)

Traduko de la Rusa poemo

“Парус” de Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27) en la Germanan de Heinrich Greif (∗1907-03-11 – †1946-07-16).

Arg-545-1944 (2015-01-24

19:29:11)

Pri Heinrich Greif vidu la retejon http: // de. wikipedia. org/ wiki/ Heinrich_Greif . La traduko troviĝas en http: // langadvice. blogspot. ru/ 2012/ 10/ blog-post. html .

Traduko de la Rusa poemo “Парус” de Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27) en la Anglan de A. Z. Foreman.

Arg-545-2269 (2015-01-28 19:03:20)

I, Manfred Retzlaff, found this poem-translation in http: // poemsintranslation. blogspot. e/ 2009/ 10/ lermontov-sail-from-russian. html .

http://www.poezio.net/version?poem-id=545&version-id=1093,1944,2269